Suppose you want to purchase a home for $525,000 with a 30-year mortgage at 4.34% interest. Suppose also that you can put down 30%. What are the monthly payments? (Round your answer to the nearest cent.)
$  

What is the total amount paid for principal and interest? (Round your answer to the nearest cent.)
$  

What is the amount saved if this home is financed for 15 years instead of for 30 years? (Round your answer to the nearest cent.)
$
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

As nothing was mentioned excel is used. If you need with formula, let me know, will do that also. Thank you.
Answer #2

1. Monthly Payment (30-Year Mortgage)

Use the mortgage formula:

M=Pr(1+r)n(1+r)n1

For 30 years (n = 360):

M=367,5000.003617(1+0.003617)360(1+0.003617)3601M367,5000.003617×3.7343.7341=367,5000.01352.734367,5000.004937$1,814.86

Monthly Payment (30-year): $1,814.86


2. Total Amount Paid (30-Year Mortgage)

Total=M×n=1,814.86×360=$653,349.60

Total Paid (30-year): $653,349.60


3. Monthly Payment (15-Year Mortgage)

For 15 years (n = 180):

M=367,5000.003617(1+0.003617)180(1+0.003617)1801M367,5000.003617×1.9351.9351=367,5000.00700.935367,5000.00749$2,752.58

Monthly Payment (15-year): $2,752.58


4. Total Amount Paid (15-Year Mortgage)

Total=2,752.58×180=$495,464.40

Total Paid (15-year): $495,464.40


5. Amount Saved (15 vs. 30 Years)

Savings=653,349.60495,464.40=$157,885.20

Amount Saved: $157,885.20



answered by: anonymous